OpenStudy (anonymous):

Which equation is in point-slope form and depicts the equation of this line? (the slopes are (2,1)(-1,-4)

OpenStudy (sleepyhead314):

first find the slope between the points, can you do that? :)

OpenStudy (anonymous):

no :(

OpenStudy (sleepyhead314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

\[y = \frac{ 5x }{ 3 } - \frac{ 7 }{ 3 }\]There is your solution to your question.

OpenStudy (anonymous):

:) i really do think yall. Yall are a BIG help.1

OpenStudy (sleepyhead314):

@Rachella in point slope form

OpenStudy (anonymous):

My pleasure! If you ever need my assistance you can tag me in your question by typing @Rachella and I'll be more than happy to help you.

OpenStudy (anonymous):

That is in point-slope form.

OpenStudy (sleepyhead314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) slope-intercept form: y = mx + b when m = slope and b = y-intercept

OpenStudy (anonymous):

(2,1) and (-1,-4) Point-Slope Form: y – y1 = m(x – x1) (2,1) and (-1,-4) in Point-Slope Form would be: \[y = \frac{ 5x }{ 3 } - \frac{ 7 }{ 3 }\]Hold on, I can give a calculator to prove myself.

OpenStudy (anonymous):

http://prntscr.com/5c7id5 Now, explain how I was wrong. @sleepyhead314